OpenStudy (anonymous):

A kindergarten school has square mats in two colors, red and blue. The ratio of the perimeter of the red mat to the perimeter of the blue mat is 2 : 1. The area of the red mat is 64 square feet. What is the area of the blue mat? 16 square feet 32 square feet 60 square feet 128 square feet

OpenStudy (anonymous):

16 square feet. Ratio of re/blue perimeter(4r_red/4r_blue)=2/1 => r_red/r_blue=2 => area of red/area of blue =4 => area of blue =(1/4)*area of red =16 sq feet